# Investigating the importance of social vulnerability in opioid-related mortality across the United States

### v2

@@ Abstract
The opioid crisis remains a critical public health challenge in the United States. Despite national efforts to reduce opioid prescribing rates by nearly 45% between 2011 and 2021, opioid overdose deaths more than tripled during this same period. This alarming trend reflects a major shift in the crisis, with illegal opioids now driving the majority of overdose deaths instead of prescription opioids. Although much attention has been given to supply-side factors fueling this transition, the underlying socioeconomic conditions that perpetuate and exacerbate opioid misuse remain less understood. Moreover, the COVID-19 pandemic intensified the opioid crisis through widespread social isolation and record-high unemployment; consequently, understanding the socioeconomic drivers of this epidemic has become even more crucial in recent years. To address this need, our study examines the correlation between opioid-related mortality and thirteen components of the Social Vulnerability Index (SVI). Leveraging a nationwide county-level dataset spanning consecutive years from 2010 to 2022, this study integrates empirical insights from exploratory data analysis with feature importance metrics derived from machine learning models. Our findings highlight critical social factors strongly correlated with opioid-related mortality, emphasizing their potential roles in worsening the epidemic when their levels are high and mitigating it when their levels are low.
@@
Significance Statement:
The opioid crisis remains one of the most pressing public health challenges in the United States, with overdose deaths more than tripling between 2011 and 2021 despite significant reductions in opioid prescribing. This study addresses a critical gap in understanding the socioeconomic factors that exacerbate or mitigate the crisis by leveraging a nationwide county-level dataset spanning over a decade. Integrating data analysis with machine learning, this study identifies key social factors that are strongly correlated with opioid-related mortality. Our findings provide actionable insights to inform interventions aimed at reducing opioid-related mortality across the nation.
Significance statement word count: 94
Introduction
@@
These unsettling trends highlight a critical transformation in the opioid epidemic: Misuse is no longer primarily sustained through prescription medications but has shifted to illicit opioids such as fentanyl, methamphetamine, and heroin [6, 7]. This transition has been driven by several factors, including the increased availability of low-cost, high-potency synthetic opioids [8, 9], the widespread mixing of drugs with unpredictable potency [10, 11, 12], the proliferation of online platforms enabling the distribution of illegal substances [13, 14], and worsening socioeconomic conditions linked to “deaths of despair” [15, 16, 17]. Although these factors had been exacerbating the epidemic well before 2020, the social isolation and record-high unemployment levels caused by the COVID-19 pandemic have further intensified the crisis [18, 19, 20, 21, 22].
@@
While much attention has been given to the supply-side factors driving the opioid crisis, the underlying socioeconomic conditions that perpetuate and worsen opioid misuse are not as well understood. A growing body of research suggests that elements of the Social Vulnerability Index (SVI)—a composite measure designed to capture a community’s resilience to external stressors—play critical roles in shaping opioid-related outcomes [23, 24, 25]. For instance, social vulnerability factors have been identified as reliable predictors of opioid-related mortality [26], counties with higher social vulnerability rates tend to exhibit greater rates of opioid use disorder (OUD) [27], and communities with greater vulnerability due to housing type or lack of transportation have reduced access to medications for OUD [25].
@@
While previous studies have provided important insights into the relationship between social vulnerability and opioid-related mortality, significant limitations persist. Many rely on cross-sectional data that does not capture trends over consecutive years and focus on aggregate SVI measures rather than its individual components. To address these gaps, this study utilizes a nationwide county-level dataset spanning consecutive years from 2010 to 2022 to examine the correlations between thirteen individual components of the SVI and opioid-related mortality. Mortality data were sourced from the CDC WONDER database, which included many suppressed values. To overcome this limitation, we developed an effective imputation method to ensure a more complete and accurate analysis of the relationships between SVI components and opioid-related mortality.
@@
To identify which SVI components are most strongly correlated with opioid-related mortality, we leverage insights from exploratory data analysis alongside feature importance results derived from machine learning models. While machine learning excels at identifying predictive variables, its results can sometimes highlight relationships that are not meaningful in real-world contexts. Conversely, data analysis uncovers empirical relationships but may fall short in predictive ability. Integrating these two methods allows us to address the limitations of each and gain a more comprehensive understanding.
@@
We begin this study by examining how the rates of each SVI component manifest in counties with anomalously high and low mortality rates. By identifying variables that exhibit their highest levels in high-mortality counties and their lowest levels in low-mortality counties, we highlight factors with bidirectional impacts across the mortality spectrum. Building on these empirical results, we then identify which SVI components are key contributors to opioid-related mortality predictions using feature importance metrics from two machine learning models.
@@
The first model, XGBoost, is a decision tree-based algorithm with a proven track record in classification and regression tasks across various domains [28, 29, 30, 31]. It evaluates feature importance using information gain, which measures how much each feature improves predictive accuracy when splitting a decision tree node. The second model, an augmented autoencoder, compresses input data into a lower-dimensional latent space and reconstructs it to predict mortality rates [32, 33, 34]. To tailor the architecture for this study, the autoencoder was augmented with a preliminary convolutional layer to aggregate the SVI components. To quantify each feature’s importance in the autoencoder, we utilized a Shapley Gradient Explainer, which identifies the variables that most significantly influence predictions when perturbed [35, 36, 37].
@@
By integrating these complementary approaches, this study advances the understanding of the complex relationships between social vulnerability and opioid-related mortality. Our findings reveal critical socioeconomic factors that are strongly correlated with opioid-related mortality, highlighting variables that may exacerbate the epidemic at high levels and mitigate it at low levels. These insights offer a valuable foundation for designing public health interventions aimed at addressing the ongoing opioid crisis.
Results
Data analysis
We sourced mortality rates from the Centers for Disease Control and Prevention’s Wide-ranging ONline Data for Epidemiologic Research (CDC WONDER) system [38], which suppresses death counts in counties with fewer than 10 deaths. Consequently, mortality rates were unavailable for nearly half of the nation’s counties in each year of the study period. To address this data censoring, we developed an effective methodology to impute missing values based on data from neighboring counties, as detailed in the Materials and Methods section.
Using this imputed dataset, our study begins by analyzing how the rates of each SVI component manifest in counties with anomalous mortality rates. In brief, each year we define anomalies as rates that fall within specific probabilities of a statistical distribution fitted to the nonzero mortality rates. Zero values were excluded because they make up a significant portion of the data, up to 947 counties, or 30.12% of the data in some years; they therefore do not represent anomalous cases. Consequently, the set of possible fitted distributions is restricted to those defined over the range $(0,\infty)$.
To determine the best-fitting distribution, we evaluated Gamma, Weibull, inverse Gaussian, and lognormal distributions. The goodness-of-fit was assessed using the Kolmogorov-Smirnov test, Akaike Information Criterion, and Bayesian Information Criterion. Based on these statistical measures, the lognormal distribution consistently outperformed the other candidates, scoring as the best-fitting distribution each and every year in the study period. Using these fitted distributions, anomalies are defined annually in two categories as follows: “Hot anomalies” are counties with nonzero rates above the 98th percentile of the yearly fitted distribution, while “cold anomalies” are those whose nonzero rates fall below the 2nd percentile. To select the size of these tails, we assessed 1%, 2%, and 3% thresholds. In some years, a 1% tail produced no cold anomalies, whereas the 2% tail always provided them. While a 3% tail also could have been used, to maximize the anomalous nature of the rates, we opted for the tightest bounds that still yielded cold anomalies for analysis each year. Figure 1 shows how the anomalies in 2016 change with 1%, 2%, and 3% tails.
To compare the rates of each SVI component within the anomalous counties, we rank the variables based on their mean rates. Specifically, we first calculate the mean rate within each year and then average these yearly means over the entire study period. Features are ordered from highest to lowest average mean rate in hot counties and from lowest to highest in cold counties. These rankings identify the variables most strongly correlated with anomalous mortality rates. The results are presented in figure 2, where yearly mean rates for each feature are shown using colored bars, and black bars indicate each variable’s average mean rate across all years of the study.
Figure 2a ranks the SVI variables based on their average mean rates in the hot counties. Among these hot counties, the component with the highest average mean rate is no vehicle (75.40), followed by below poverty (75.22), no high school diploma (73.48), unemployment (70.18), and mobile homes (68.61). The high ranking of these features means they tend to manifest their highest rates in counties with anomalously high opioid-related mortality. This suggests that opioid-related mortality may be exacerbated by high levels of these factors. Beyond these top five, rates noticeably decline, starting with single-parent households (54.61) and continuing to drop further down the list. At the bottom are limited English ability and minority status, meaning they have the lowest average rates in the hot counties. This suggests that these components may contribute the least to the exceptionally high mortality rates observed in the hot counties.
Figure 2b ranks the SVI variables based on their average mean rates in the cold counties. Among these cold counties, the component with the lowest average mean rate is unemployment (35.58), followed by multi-unit structures (39.66), single-parent households (39.89), limited English ability (42.22), and no vehicle (42.84). The high ranking of these features means they tend to manifest their lowest rates in counties with anomalously low opioid-related mortality. This suggests that opioid-related mortality may be alleviated by lower levels of these factors. Beyond these top five, mean rates only gradually increase, except at the bottom, where a significant jump is observed for the variable aged 65 or older. This factor stands out with the highest average mean rate (61.74) in the cold anomalies, indicating that counties with anomalously low opioid-related mortality tend to have a significantly higher population of persons aged 65 or older.
Notably, only two components—unemployment and lack of vehicle access—exhibited bidirectional relationships, with their highest rates occurring in counties with anomalously high mortality and their lowest rates in counties with anomalously low mortality. The fact that these variables are observed at both extreme ends of the mortality rate spectrum suggests that they may play a particularly significant role: High levels may contribute to increased opioid-related mortality, while low levels may help reduce it. These empirical results provide a real-world foundation for the subsequent machine learning analysis, identifying which SVI features are the key contributors to opioid-related mortality predictions.
Efficacy of machine learning models
This study employed two machine learning models to predict opioid-related mortality rates from 2011 to 2022 using the previous year’s SVI data. XGBoost generates predictions for each county by combining the outputs from an ensemble of decision trees, where each tree iteratively corrects the errors of its predecessors through gradient-based optimization. The autoencoder model, on the other hand, processes input data through an initial 1D convolutional layer that aggregates SVI variables before passing the data to an encoder, which compresses it into a lower-dimensional representation. This representation is then decoded to produce the yearly mortality rate predictions.
The efficacy of our machine learning models in forecasting county-level mortality rates is evaluated using both accuracy and error metrics. For each county, the error is measured as the absolute residual. To assess overall accuracy, the maximum absolute residual, which indicates the largest annual prediction error, is used as a benchmark: Each county’s error is normalized by this highest observed error across the country. The performance of the models is summarized in table 2, which presents the average errors, maximum errors, and average accuracy of both models for each predictive year in the study period. While each year is treated as an independent test year by XGBoost, 2022 serves as the sole testing year for the autoencoder. This approach was chosen to preserve the natural temporal trajectory of the opioid crisis by training the autoencoder sequentially on data from 2010 to 2020, with 2021 as the testing year and predictions made for 2022. Accordingly, we present visual representations of the 2022 efficacy results using error histograms (fig. 3) and accuracy maps (fig. 4).
XGBoost feature importance
We used XGBoost to predict opioid-related mortality rates from 2011 to 2022 based on the previous year’s SVI data. To assess the importance of each SVI component, XGboost utilizes information gain:
This metric, detailed in the materials and methods section, quantifies the improvement in the model’s predictive performance when a feature is used to split a node in a decision tree. As a result, the features with the highest gain scores are those that most significantly improve the model’s ability to predict opioid-related mortality rates.
Autoencoder feature importance
In this study, the autoencoder model was used to predict opioid-related mortality rates from 2011 to 2022 based on the previous year’s SVI data. To determine which components of the SVI were most important in the autoencoder’s predictions, we applied a Shapley Gradient Explainer to the initial convolutional layer. The Shapley Gradient Explainer generates SHAP (Shapley Additive Explanations) values, which quantify the contribution of each feature to a model’s predictions [35, 36, 37]. By measuring how sensitive the autoencoder’s mortality predictions are to small changes in each feature, the SHAP values identify the most important components as those for which minor changes lead to the greatest impact on predictions.
The variable with the highest average SHAP value is crowding (0.00143), followed closely by no vehicle (0.00137) and multi-unit structures (0.00136). In fourth place is aged 65 or older (0.00131), from which SHAP values gradually decrease down to unemployment (0.00116) in seventh place. Beyond unemployment, there is a noticeable drop in the SHAP values for mobile homes and minority status. After which, there is another steep decline in importance moving down to the the last four components. This reflects the much smaller contribution these four factors make to the model’s predictions. Notably, the high importance score of no vehicle corroborates the empirical findings from the data analysis.
SHAP values are relative to the model, data, and problem context; they reflect the contributions of each feature to the predictions compared to the baseline (i.e., the average prediction). Ranking these values within the model provides a reliable and interpretable measure of feature importance, offering meaningful insights into the influence of each component on this specific model and dataset.
Discussion
Utilizing thirteen components of the SVI, this study sought to identify which variables are most strongly correlated with opioid-related mortality. The data analysis portion of this study examined how the rates of each component manifested in counties with anomalous mortality. Notably only two factors—unemployment and lack of vehicle access—exhibited bidirectional relationships, with their highest rates observed in counties with anomalously high mortality and their lowest rates in counties with anomalously low mortality. This pattern suggests that reducing unemployment and improving vehicle access may lead to measurable reductions in opioid-related mortality, particularly in the most severely affected regions.
These empirical results provided real-world grounding for the subsequent machine learning analysis, which identified the SVI components that are key contributors to opioid-related mortality predictions. This portion of the study employed two machine learning models: XGBoost and an autoencoder. Feature importance was assessed using information gain for XGBoost and a Shapley gradient explainer for the autoencoder. While the feature importance rankings differed between the two models, this difference highlights the value of employing multiple approaches, as it revealed critical insights that might have been overlooked if only one model had been used.
Public health interventions could mitigate these impacts by prioritizing strategies that address both economic stability and transportation access. Workforce development programs that focus on providing stable, meaningful employment opportunities could play a critical role in reducing the socioeconomic stressors that contribute to opioid misuse, particularly in regions with high opioid-related mortality. Additionally, improving access to reliable and affordable transportation, especially in underserved communities and rural regions, could help alleviate the barriers faced by those without personal vehicles. By addressing these fundamental social vulnerabilities, targeted interventions could help reduce the impact of these two critical factors and may even promote long-term recovery in the nation’s most vulnerable areas.
This study has important limitations to note, beginning with the issue of missing mortality data in the CDC WONDER dataset. Although we employed methods to address this, no approach can fully overcome the challenges posed by suppressed data. This limitation was sometimes evident when real data, missing in the previous year, became available the following year at a much higher rate than the previously imputed value. This sharp increase sometimes led to large predictive errors in both models. Future research could apply the methodologies outlined in this study to a dataset with fewer missing values, if such data is available.
The next limitation to note is that this study does not stratify counties by urban or rural designation; there are stark differences in how vehicle access impacts opioid-related mortality within these two contexts. In urban areas, such as New York City, low vehicle ownership rates are often offset by robust public transportation systems, making personal vehicles less essential for accessing healthcare. Conversely, in rural regions, such as those in Appalachia, lacking a vehicle often equates to extreme isolation, severely limiting access to medical services and treatment. Future studies could explore such urban-rural dynamics to develop context-specific strategies that address transportation barriers in a manner tailored to local needs.
Finally, the opioid crisis is a highly complex issue, influenced by many factors beyond those considered in this study. Future work could expand on the methodologies used here by investigating a broader range of factors contributing to opioid-related mortality.
Materials and Methods
Mortality data
@@ Discussion
Method for handling missing data
@@
To demonstrate the effectiveness of our methodology, we present national mortality rate maps for 2013. This year had the most significant data censoring, with 1,681 out of 3,144 counties lacking mortality rates, or 53.47% of the data. Figure 6 displays maps both with and without the missing data, each colored by percentiles from the respective empirical distributions. Figure 6a illustrates the extent of the missing data, revealing an incomplete view of national mortality rates. In contrast, Figure 6b shows the results after applying our imputation method; previously missing regions are now filled with interpolated data that preserves the spatial patterns of the original dataset. This significant improvement highlights the effectiveness of our approach, allowing us to address substantial data gaps and create a more complete and accurate national dataset for analysis.
@@
Figure 7 demonstrates the results of using the Tobler library to interpolate mortality rates for Connecticut in 2021, the last year before the county structure changed. Figure 7a displays Connecticut’s old county structure (pre-2022), while Figure 7b shows the new 2022 boundaries. For direct comparison, the color scales for both maps are normalized to the minimum and maximum 2021 mortality rates across both county structures. As an example, counties 09003 and 09013 from the old structure roughly correspond to county 09110 in the new 2022 structure. Here the interpolated data accurately reflects this consolidation, with the interpolated value being substantially larger than both old values. For another example, county 09009 in the old structure roughly corresponds to counties 09140 and 09170 in the new structure. In this example, the interpolated data captures this division effectively, with both interpolated values considerably lower than the old value.
SVI data
This study uses county-level rank ordered percentile estimates (ranging from 0 to 100) of thirteen components of the SVI across the United States from 2010 to 2022. The set of SVI variables utilized for 2010 data collection differs from those used in 2022. Therefore, we were forced to exclude all variables that were not available across every year in the study period, resulting in the thirteen SVI features utilized in this study. All data were sourced from the CDC and the Agency for Toxic Substances and Disease Registry’s Social Vulnerability Index, which utilizes data from the American Community Survey to assess the resilience of communities to external stresses on human health [23].
The 2010 SVI data was not followed by another release until 2014, after which the data has been released biennially. To account for the gap between 2010 and 2014, we imputed the intervening years by calculating a quarter of the difference between the 2010 and 2014 data points and adding it consecutively to each year starting in 2014. After 2014, for the biennial data releases, we applied a similar method by calculating half the difference between consecutive data points and adding it to the earlier year’s rates. This method, with the understanding that yearly rate fluctuations are modest, provided a viable way to analyze the impact of these social vulnerability factors on the national opioid crisis throughout the entire study period.

### v3

# Investigating the importance of county-level characteristics in opioid-related mortality across the United States

@@ Abstract
The opioid crisis remains a critical public health challenge in the United States. Despite national efforts which reduced opioid prescribing rates by nearly 45% between 2011 and 2021, opioid-related overdose deaths more than tripled during the same period. This alarming trend reflects a major shift in the crisis, with illegal opioids now driving the majority of overdose deaths instead of prescription opioids. Although much attention has been given to supply-side factors fueling this transition, the underlying structural conditions that perpetuate and exacerbate opioid misuse remain less understood. Moreover, the COVID-19 pandemic intensified the opioid crisis through widespread social isolation and record-high unemployment; consequently, understanding the underlying drivers of this epidemic has become even more crucial in recent years. To address this need, our study examines the correlation between opioid-related mortality and thirteen county-level characteristics related to population traits, economic stability, and infrastructure. Leveraging a nationwide county-level dataset spanning consecutive years from 2010 to 2022, this study integrates empirical insights from exploratory data analysis with feature importance metrics derived from machine learning models. Our findings highlight critical regional characteristics strongly correlated with opioid-related mortality, emphasizing their potential roles in worsening the epidemic when their levels are high and mitigating it when their levels are low.
@@
1  Introduction
@@
These unsettling trends highlight a critical transformation in the opioid epidemic: Misuse is no longer primarily sustained through prescription medications but has shifted to illicit opioids such as fentanyl or heroin [6, 7]. This transition has been driven by several factors, including the increased availability of low-cost, high-potency synthetic opioids [8, 9], the widespread mixing of drugs with unpredictable potency [10, 11, 12], the proliferation of online platforms enabling the distribution of illegal substances [13, 14], and worsening regional conditions associated with preventable mortality linked to mental health and substance use[15, 16, 17]. Although these factors had been exacerbating the epidemic well before 2020, the social isolation and record-high unemployment levels caused by the COVID-19 pandemic have further intensified the crisis [18, 19, 20, 21, 22].
@@
While much attention has been given to the supply-side factors driving the opioid crisis, the underlying structural conditions that perpetuate and exacerbate opioid misuse are not as well understood. A growing body of research suggests that elements of the Social Vulnerability Index (SVI)—a composite measure designed to capture a community’s resilience to external stressors—play critical roles in shaping opioid-related outcomes [23, 24, 25]. Defined in table 1, the features in the SVI are county-level characteristics that have been identified as reliable predictors of opioid-related mortality [26]; counties with elevated levels in these measures tend to exhibit higher rates of opioid use disorder (OUD) [27], and communities with worse housing conditions and less access to transportation often face reduced access to medications for OUD [25].
Theme Feature Definition
Socioeconomic Status Below poverty Percentile percentage of persons below poverty estimate
Unemployment Percentile percentage of civilian (age 16+) unemployed estimate
No high school diploma Percentile percentage of persons with no high school diploma (age 25+) estimate
Household Characteristics Age 65 and older Percentile percentage of persons aged 65 and older estimate
Age 17 and younger Percentile percentage of persons aged 17 and younger estimate
Single-parent households Percentile percentage of single-parent households with children under 18 estimate
Limited English ability Percentile percentage of persons (age 5+) who speak English “less than well” estimate
Racial and Ethnic Minority Status Minority Status Percentile percentage minority (Hispanic or Latino (of any race); Black and African American, Not Hispanic or Latino; American Indian and Alaska Native, Not Hispanic or Latino; Asian, Not Hispanic or Latino; Native Hawaiian and Other Pacific Islander, Not Hispanic or Latino; Two or More Races, Not Hispanic or Latino; Other Races, Not Hispanic or Latino) estimate
Housing Type and Transportation Multi-unit structures Percentile percentage of housing structures with 10 or more units estimate
Mobile homes Percentile percentage mobile homes estimate
Crowding Percentile percentage of households with more people than rooms estimate
No vehicle Percentile percentage of households with no vehicle available estimate
Group quarters Percentile percentage of persons in group quarters estimate (e.g., correctional institutions, nursing homes)
Table 1: Definitions of the county-level characteristics utilized in this study.
@@
While previous studies have provided important insights into the relationship between these county-level characteristics and opioid-related mortality, significant limitations persist. Many rely on cross-sectional data that does not capture trends over consecutive years and focus on aggregate measures rather than individual components. To address these gaps, this study utilizes a nationwide county-level dataset spanning consecutive years from 2010 to 2022 to examine the correlations between thirteen individual county-level characteristics and opioid-related mortality. Mortality data were sourced from the CDC WONDER database, which included many suppressed values. To overcome this limitation, we developed an effective imputation method to ensure a more complete and accurate analysis of the relationships between these area-based measures and opioid-related mortality.
@@
To identify which county-level characteristics are most strongly correlated with opioid-related mortality, we leverage insights from exploratory data analysis alongside feature importance results derived from machine learning models. While machine learning excels at identifying predictive variables, its results can sometimes highlight relationships that are not meaningful in real-world contexts. Conversely, data analysis uncovers empirical relationships but may fall short in predictive ability. Integrating these two methods allows us to address the limitations of each and gain a more comprehensive understanding.
While traditional regression models are well-suited for estimating marginal effect sizes under assumptions of linearity and additivity, they are limited in capturing complex, nonlinear interactions that may drive opioid-related mortality. In contrast, machine learning approaches can model these complexities without prespecifying a functional form. This study prioritizes the latter capability, as our goal is to uncover potentially nonlinear patterns across diverse county contexts rather than restrict our analysis to relationships that can be expressed in a simple linear framework. Additionally, machine learning and artificial intelligence approaches have been successfully applied in medicine to detect patterns and uncover relationships [28, 29, 30, 31], further motivating their use here.
@@
We begin this study by examining how the rates of each characteristic manifest in counties with anomalously high and low mortality rates. By identifying variables that exhibit their highest levels in high-mortality counties and their lowest levels in low-mortality counties, we highlight factors with bidirectional impacts across the mortality spectrum. This exploratory data analysis provides interpretable, preliminary insights into variable relationships before applying machine learning methods. We then identify which characteristics are key contributors to opioid-related mortality predictions using feature importance metrics from two machine learning models.
@@
The first model, XGBoost, is a decision tree-based algorithm with a proven track record in classification and regression tasks across various domains [32, 33, 34, 35]. It evaluates feature importance using information gain, which measures how much each feature improves predictive accuracy when splitting a decision tree node. The second model, an augmented autoencoder, compresses input data into a lower-dimensional latent space then decodes it back into the county-level space to produce mortality rate predictions [36, 37, 38]. To tailor the architecture for this study, the autoencoder was augmented with a preliminary convolutional layer to aggregate the thirteen SVI variables. To quantify each feature’s importance in the autoencoder, we utilized a Shapley Gradient Explainer, which identifies the variables that most significantly influence predictions when perturbed [39, 40, 41].
@@
By integrating these complementary approaches, this study advances the understanding of the complex relationships between county-level characteristics and opioid-related mortality. Our findings reveal critical structural factors that are strongly correlated with opioid-related mortality, highlighting variables that may exacerbate the epidemic at high levels and mitigate it at low levels. These insights offer a valuable foundation for designing public health interventions aimed at addressing the ongoing opioid crisis.
2  Materials and methods
2.1  Mortality data
@@
2.2  Method for handling missing data
@@
To evaluate the effectiveness of our imputation strategy, we conducted an efficacy comparison using a complete national-level dataset, spanning 2014 to 2020, of opioid-related mortality rates sourced from HepVu [47]. To simulate the level of missingness present in our study data, we randomly censored approximately half the county rates in each year. This allowed us to assess the efficacy of various imputation techniques by comparing each method’s imputed values to the original, uncensored rates.
We compared our method against three alternative strategies. The first was national mean imputation, where missing values were imputed using the national average of all available county rates. The second was state mean imputation [46], where missing values were imputed using the average of available county rates within the same state. This approach suffered, however, in years where entire states had no uncensored counties, in which case we defaulted to the national mean. The third was inverse distance weighting (IDW), where for each missing county, available neighbors were weighted by the inverse of their geodesic distance from the target county’s centroid. Widely used in hydrology for estimating rainfall [48, 49], IDW assigns greater influence to nearby locations and is a long-established method for geospatial interpolation.
The results of this comparison are presented in table 2, with additional histograms available in the project’s GitHub repository. Performance was evaluated using mean absolute error (MAE), root mean squared error (RMSE), mean percentage error (MPE), and mean absolute percentage error (MAPE). Across all years, our method substantially outperformed both the national and state-level imputations. As seen in table 2, while the IDW method achieved marginally lower errors in some years, our novel approach performed nearly identically. Beyond its strong performance, our neighbor-based method integrates more naturally into machine learning and deep learning workflows, yielding reproducible feature values without the added complexity of distance calculations or parameter tuning. Additionally, its simplicity, interpretability, and near-parity with IDW make it a highly practical choice; particularly in geospatial contexts where adjacency is both intuitive and meaningful.
Year Method MAE RMSE MPE MAPE
2014 National mean 5.67 7.71 -34.09% 55.48%
State mean 4.37 6.30 -16.19% 35.03%
IDW weighted mean 3.35 4.83 -9.21% 25.87%
Adjacent neighbors mean 3.38 4.91 -9.70% 26.14%
2015 National mean 6.23 8.40 -30.53% 54.12%
State mean 4.63 6.56 -13.70% 33.92%
IDW weighted mean 3.69 5.28 -7.66% 25.54%
Adjacent neighbors mean 3.71 5.28 -8.03% 25.72%
2016 National mean 7.54 10.24 -35.14% 57.98%
State mean 5.34 7.73 -15.58% 34.45%
IDW weighted mean 4.21 6.41 -7.71% 24.93%
Adjacent neighbors mean 4.20 6.38 -7.91% 24.98%
2017 National mean 8.20 11.17 -35.87% 58.90%
State mean 5.73 8.41 -14.71% 34.20%
IDW weighted mean 4.54 6.78 -8.11% 25.63%
Adjacent neighbors mean 4.53 6.78 -8.19% 25.67%
2018 National mean 7.83 10.43 -36.94% 59.65%
State mean 5.46 7.82 -15.26% 34.20%
IDW weighted mean 4.33 6.26 -9.03% 25.90%
Adjacent neighbors mean 4.33 6.26 -9.49% 26.08%
2019 National mean 7.82 10.97 -30.29% 54.67%
State mean 5.84 8.61 -14.37% 34.62%
IDW weighted mean 4.59 7.03 -5.73% 25.23%
Adjacent neighbors mean 4.60 7.00 -6.00% 25.35%
2020 National mean 10.37 15.01 -32.20% 55.67%
State mean 7.60 11.65 -13.08% 34.01%
IDW weighted mean 6.08 9.55 -5.45% 25.49%
Adjacent neighbors mean 6.09 9.56 -5.70% 25.58%
Table 2: Performance comparison of four mortality rate imputation methods. The global and state mean methods use national and state-level averages, respectively. The IDW weighted method imputes rates using inverse distance weighting of nearby counties, while the adjacent neighbors method (our novel approach) imputes using the average of available queen-adjacent neighbors. Metrics include mean absolute error (MAE), root mean squared error (RMSE), mean percentage error (MPE), and mean absolute percentage error (MAPE).
@@
To showcase the results of our methodology, we present national mortality rate maps for 2013. This year had the most significant data censoring, with 1,681 out of 3,144 counties lacking mortality rates, or 53.47% of the data. Figure 1 displays maps both with and without the missing data, each colored by percentiles from the respective empirical distributions. Figure 1a illustrates the extent of the missing data, revealing an incomplete view of national mortality rates. In contrast, fig. 1b shows the results after applying our imputation method; previously missing regions are now filled with interpolated data that preserves the spatial patterns of the original dataset. This significant improvement highlights the effectiveness of our approach, allowing us to address substantial data gaps and create a more complete and accurate national dataset for analysis.
Figure 1: Missing and interpolated data maps for the 2013 mortality rates. (A) 2013 missing data map. (B) 2013 interpolated data map. Each map is colored by percentiles from the respective empirical distributions. Figures are best viewed in color.
@@
Figure 2 demonstrates the results of using the Tobler library to interpolate mortality rates for Connecticut in 2021, the last year before the county structure changed. Figure 2a displays Connecticut’s old county structure (pre-2022), while fig. 2b shows the new 2022 boundaries. For direct comparison, the color scales for both maps are normalized to the minimum and maximum 2021 mortality rates across both county structures. As an example, counties 09003 and 09013 from the old structure roughly correspond to county 09110 in the new 2022 structure. Here the interpolated data accurately reflects this consolidation, with the interpolated value being substantially larger than both old values. For another example, county 09009 in the old structure roughly corresponds to counties 09140 and 09170 in the new structure. In this example, the interpolated data captures this division effectively, with both interpolated values considerably lower than the old value.
Figure 2: Comparison of the 2021 mortality rates using the old and new Connecticut county structures. (A) 2021 mortality rates for the old (pre-2022) Connecticut county structure. (B) 2021 mortality rates for the new (2022) Connecticut county structure. The color scales for both maps are normalized to the minimum and maximum 2021 mortality rates across both county structures. The minimum 2021 mortality rate is 6.09, shown in the darkest blue, while the maximum rate is 68.45, shown in the darkest red. Figures are best viewed in color.
2.3  County-level characteristic data
This study uses county-level rank ordered percentile estimates (ranging from 0 to 100) of thirteen SVI variables across the United States from 2010 to 2022. The set of variables utilized for 2010 data collection differs from those used in 2022. Therefore, we were forced to exclude all factors that were not available across every year in the study period, resulting in the thirteen characteristics utilized in this study. All data were sourced from the CDC and the Agency for Toxic Substances and Disease Registry’s Social Vulnerability Index, which utilizes data from the American Community Survey to assess the resilience of communities to external stresses on human health [23].
The 2010 data was not followed by another release until 2014, after which the data has been released biennially. To account for the gap between 2010 and 2014, we imputed the intervening years by calculating a quarter of the difference between the 2010 and 2014 data points and adding it consecutively to each year starting in 2014. After 2014, for the biennial data releases, we applied a similar method by calculating half the difference between consecutive data points and adding it to the earlier year’s rates. This method, with the understanding that yearly rate fluctuations are modest, provided a viable way to analyze the impact of these county-level characteristics on the national opioid crisis throughout the entire study period.